# Existence and Uniqueness Results for a Mean-field Game of Optimal Investment

### v2

@@ 4.2. The infinite time horizon case
and hence $\widehat{q}$ verifies condition (4.3).
Case $x>y_{\infty}$. From Theorem 4.12, we know that $\widehat{q}$ is monotone decreasing and converges to $y_{\infty}$. This implies that, $0<y_{\infty}<\widehat{q}_{s}\leq x$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by the constant $K$ defined in (4.37). Therefore, $\widehat{q}$ satisfies condition (4.3).
Case $x<y_{\infty}$. From Theorem 4.12, we know that $\widehat{q}$ is monotone increasing and converges to $y_{\infty}$. This implies that, $0<x\leq\widehat{q}_{s}<y_{\infty}$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by $\frac{x^{-\beta}}{\rho+\delta}$. Therefore, $\widehat{q}$ satisfies condition (4.3).
@@
Also in this case the results of this section can be easily extended to the case where we consider an initial time time $t>0$.
@@
$\begin{dcases}\dfrac{{\mathrm{d}}}{{\mathrm{d}}s}y_{s}=-\delta y_{s}+\int_{s}^{T}{\mathrm{e}}^{-(\rho+\delta)(u-s)}y_{u}^{-\beta}\,{\mathrm{d}}u,&s\in[t,T],\\ y_{t}=x\,,\end{dcases}$ (4.38)
@@
The statements of Theorems 4.9 and 4.12, and of Propositions 4.10 and 4.13, remain the same (except for minor modifications). The bounds given in (4.14) become
@@ 5. The deterministic model
In this section we discuss the deterministic version of the mean-field game problem introduced in Section 2. The structure of the problem will allow us to exploit the results given in Sections 3 and 4, to get the unique equilibrium of the mean-field game in an explicit form.
@@
where $\delta>0$ is a given coefficient and the control ${\mathbf{u}}\coloneqq(u_{s})_{s\in[0,T]}$ is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
@@
[…]
Note that, for any $x\in\mathbb{R}$ and any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, Equation (5.1) has a unique solution $X^{x,{\mathbf{u}}}$, given by
As in the stochastic case, $X$ describes the evolution of the production capacity of a representative agent, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate ${\mathbf{u}}\in{\mathcal{U}}_{T}$.
Next, we consider the discounted net profit functional
where $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function.
@@
Since the representative agent is chosen randomly by picking her/his initial production capacity level $x$ according to the initial distribution $\nu_{0}$, the assumption above ensures that, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, $X_{s}^{x,{\mathbf{u}}}\geq 0$, for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative.
@@
$\begin{dcases}\dfrac{\partial}{\partial s}p(s,x)+\dfrac{\partial}{\partial x}\left(p(s,x)[-\delta x+u_{s}]\right)=0,&(s,x)\in(0,T]\times(0,+\infty)\,,\\ p(0,x)=p_{0}(x),&x\in(0,+\infty)\,,\end{dcases}$ (5.6)
@@
Proof.
Fix ${\mathbf{u}}\in{\mathcal{U}}_{T}$ and $p_{0}$ as above. The fact that $p^{p_{0},{\mathbf{u}}}$ is a solution to (5.6) can be verified via standard computations (see, e.g., [5, Example 2.5.2] or [7, Lemma 4.15]). Uniqueness follows from an argument similar to the one used in the proof of [7, Lemma 4.16].
@@
A pair $(\widehat{\mathbf{u}},\widehat{q})$, where $\widehat{\mathbf{u}}\in{\mathcal{U}}_{T}$ and $\widehat{q}\colon[0,T]\to(0,+\infty)$ is a measurable function, is an equilibrium of the deterministic mean-field game if
@@ 5.1. The optimization problem
In the finite time horizon case, i.e. $T<+\infty$, the following result holds, which is analogous to Theorem 3.3.
Theorem 5.5.
Fix $x>0$ and $q\colon[0,T]\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty))$. Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.4), is an optimal control for problem (5.10).
@@
Proof.
Replicating the argument of the proof of Proposition 3.1, we get that
@@
Then, the result follows using the same reasoning as in the proof of Theorem 3.3. ∎
In the infinite time horizon case, i.e. $T=+\infty$, we have the following statement, which is the deterministic counterpart of Theorem 3.6. We state it without proof.
Theorem 5.6.
Fix $x>0$ and consider a function $q\colon[0,+\infty)\to(0,+\infty)$ such that $q^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))$ and such that the function $z^{(\infty,q)}$, defined in (3.10) is bounded on $[0,+\infty)$.
Then, $\widehat{\mathbf{u}}^{(\infty,q)}\coloneqq z^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (5.10). Moreover, $\widehat{\mathbf{u}}^{(\infty,q)}$ is essentially unique, i.e., if $\overline{\mathbf{u}}^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (5.10) different from $\widehat{\mathbf{u}}^{(\infty,q)}$, then
@@ 5.2. Existence and uniqueness of equilibria
Also in the deterministic mean-field game previously introduced, the search for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 5.4, this map is $(q_{s})_{s\in[0,T]}\mapsto\left(\int_{0}^{+\infty}x\,m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}(s,x)\,{\mathrm{d}}x\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the optimization problem (5.10) – whose expression is given either in Theorem 5.5, in the case $T<+\infty$, or in Theorem 5.6, in the case $T=+\infty$ – and $m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}$ is the unique solution to (5.6), with initial probability density $m_{0}$ and control $\widehat{\mathbf{u}}^{(T,q)}$.
@@
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 5.4, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}((0,T];(0,+\infty)),$ (5.15)
    then $\widehat{q}$ is a solution to the integral equation
@@
    Vice versa, if there exist a unique solution $\widehat{q}$ to (5.16) satisfying (5.15), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(T,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (5.15), where $z^{(T,q)}$ is the function defined in (3.4).
@@
    If there exists an equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ in the sense of Definition 5.4, such that
    $\widehat{q}^{-\beta}\in{\mathrm{L}}^{1}_{\rho+\delta}((0,+\infty);(0,+\infty))\quad\text{and}\quad z^{(\infty,\widehat{q})}\text{ is bounded on }[0,+\infty)\,,$ (5.17)
    where $z^{(\infty,q)}$ is the function defined in (3.10), then $\widehat{q}$ is a solution to the integral equation
@@
    Vice versa, if there exist a unique solution $\widehat{q}$ to (5.18) satisfying (5.17), then there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})=(z^{(\infty,\widehat{q})},\widehat{q})$ of the mean-field game among all equilibria $(\widetilde{\mathbf{u}},\widetilde{q})$ such that $\widetilde{q}$ verifies (5.17).
Finally, we get the following result, which establishes existence and uniqueness of an equilibrium also for the deterministic mean-field game. We omit its proof, since it follows the same lines of the proofs of Propositions 4.10 and 4.13, except for minor adaptations.
Proposition 5.8.
For each fixed $\nu_{0}$ satisfying Assumption 5.1, there exists a unique equilibrium $(\widehat{\mathbf{u}},\widehat{q})$ for the deterministic mean-field game with finite time horizon (resp., infinite time horizon), among all equilibria verifying the integrability condition (5.15) (resp., (5.17)).

### v3

@@ 4.2. The infinite time horizon case
and hence $\widehat{q}\in{\mathcal{Q}}_{\infty}$.
Case $x>y_{\infty}$. From Theorem 4.11, we know that $\widehat{q}$ is monotone decreasing and converges to $y_{\infty}$. This implies that, $0<y_{\infty}<\widehat{q}_{s}\leq x$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by the constant $K$ defined in (4.35). Therefore, $\widehat{q}\in{\mathcal{Q}}_{\infty}$.
Case $x<y_{\infty}$. From Theorem 4.11, we know that $\widehat{q}$ is monotone increasing and converges to $y_{\infty}$. This implies that, $0<x\leq\widehat{q}_{s}<y_{\infty}$, for all $s\geq 0$, and that $z^{(\infty,\widehat{q})}$ is positive and bounded above by $\frac{x^{-\beta}}{\rho+\delta}$. Therefore, $\widehat{q}\in{\mathcal{Q}}_{\infty}$.
@@
A crucial question regarding the robustness of the mean-field equilibrium is whether the aggregate behavior of the agents remains consistent when passing from a large but finite horizon to an infinite one. Thanks to the characterization of the equilibrium provided by Theorem 4.1 and the analytical results obtained in Theorem 4.11, we can formally establish that the finite-horizon equilibrium $\widehat{q}^{(T)}$ converges to the unique infinite-horizon equilibrium $\widehat{q}^{(\infty)}$. This result, summarized in the following corollary, ensures the stability of the model’s predictions with respect to the time horizon $T$ and justifies the use of the infinite-horizon case as a structural limit for long-term strategic interactions.
Corollary 4.13 (Convergence of the Mean-Field Equilibrium).
For each $T>0$, let $(\widehat{\mathbf{u}}^{(T)},\widehat{q}^{(T)})$ be the unique equilibrium of the mean-field game with finite horizon $T$, where $\widehat{q}^{(T)}$ is the unique solution to (4.4). Under the assumptions of Theorem 4.11, as $T\to+\infty$, the sequence of equilibrium trajectories $\{\widehat{q}^{(T)}\}_{T>0}$ converges to the unique infinite-horizon equilibrium $\widehat{q}^{(\infty)}$ in the weighted space ${\mathrm{L}}^{1}_{\rho+2\delta}([0,\infty);\mathbb{R})$. Consequently, the optimal strategies $\widehat{\mathbf{u}}^{(T)}$ converge to the infinite-horizon optimal strategy $\widehat{\mathbf{u}}^{(\infty)}$ in the same sense.
Dimostrazione.
The proof of Theorem 4.11 shows that the family of transformed solutions ${\mathcal{Z}}\coloneqq\{z^{(T)}\}_{T>0}$ is relatively compact in ${\mathrm{L}}^{1}_{\rho+2\delta}$ by means of the Fréchet-Kolmogorov theorem. Let $t_{n}\to\infty$ be an arbitrary sequence, consider the corresponding sequence $\{z^{(t_{n})}\}$, and take any subsequence $\{z^{(t_{n_{k}})}\}$. By the existence part of the proof of Theorem 4.11, its limit must solve the infinite-horizon integro-differential equation. Since the uniqueness part of the same Theorem guarantees that such a solution is unique within the admissible set ${\mathcal{C}}_{x,\infty}$, we conclude that $z^{(t_{n_{k}})}\to z^{(\infty)}$. Since this holds for an arbitrary sequence $t_{n}\to\infty$, we conclude that the entire family $\mathcal{Z}$} converges to $z^{(\infty)}$ as $T\to\infty$. By the continuous mapping between the auxiliary function $z$ and the equilibrium trajectory $\widehat{q}$ (and the corresponding optimal control $\widehat{\mathbf{u}}$), the convergence of the equilibrium follows. ∎
@@
Also in this case the results of this section can be easily extended to the case where we consider an initial time $t>0$.
@@
[…]
@@
The statements of Theorems 4.8, 4.9, 4.11, and 4.12, remain the same (except for minor modifications). The bounds given in (4.12) become
@@
4.2.1. Economic Interpretation of the Stationary equilibrium
The explicit nature of the stationary solution in the infinite horizon problem allows for a clear economic characterization of the long-term market equilibrium. In this setting, the unique stationary average production level $y_{\infty}$ is the positive solution to the algebraic equation $(\rho+\delta)\delta y-y^{-\beta}=0$ (cf. (4.28)), given by:
$y_{\infty}=\left[\frac{1}{\delta(\rho+\delta)}\right]^{\frac{1}{1+\beta}}.$
Some comments on the economics of this result are in place.
Sensitivity to parameters. The term $\delta(\rho+\delta)$ can be interpreted as the effective user cost of capital. It combines the physical depreciation rate $\delta$ with the opportunity cost of capital $\rho$. As either the discount rate $\rho$ or the depreciation rate $\delta$ increases, the stationary production capacity $y_{\infty}$ decreases. This reflects the intuition that firms are less willing to maintain high production levels when the future is heavily discounted or when maintaining capacity requires more intensive reinvestment. Furthermore, the parameter $\beta$ represents the inverse of the price elasticity of demand. A higher $\beta$ (less elastic demand) leads to a lower $y_{\infty}$, as firms strategically limit capacity to avoid the sharp price decreases associated with high aggregate supply in inelastic markets.
Market stability and convergence. The monotone convergence result of Theorem 4.11 provides further economic insight. If the initial production level exceeds $y_{\infty}$, the equilibrium path decreases monotonically toward the steady state; if it starts below $y_{\infty}$, it increases monotonically. Thus, the stationary equilibrium is globally stable and dynamically attractive. Deviations from the steady state are gradually absorbed through investment adjustments without oscillations or overshooting, suggesting that competitive investment incentives dampen imbalances rather than amplify them. Notably, the stationary level $y_{\infty}$ and the trajectory of the average production $\widehat{q}$ are independent of the volatility parameter $\sigma$. This suggests that while individual firms face stochastic fluctuations, the "mean-field" or aggregate economy follows a deterministic and stable path toward equilibrium. We refer also to Section 5 below on this.
@@ 5. The deterministic model
In this section we discuss the deterministic version of the mean-field game problem introduced in Section 2. The structure of the problem will allow us to exploit the results given in Sections 3 and 4, to get the unique equilibrium of the mean-field game in an explicit form. Indeed, the results obtained for the stochastic model reveal that the average production capacity and the optimal control at equilibrium do not depend on parameter $\sigma$ appearing in SDE (2.1), which describes the sensitivity of the production capacity with respect to random shocks. Thus, by formulating a deterministic counterpart of the problem studied in the previous sections, i.e., formally setting $\sigma=0$ in the dynamics of the production capacity, we are able to show that results similar to those proved in Sections 3 and 4 are valid also in this context.
@@
where $\delta>0$ is a given coefficient and ${\mathbf{u}}\coloneqq(u_{s})_{s\in[0,T]}\in{\mathrm{L}}^{1}([0,T];[0,+\infty))$ is the control, which we will choose in a suitable class of admissible functions, that will be specified below.
Note that, for any $x\in\mathbb{R}$ and any ${\mathbf{u}}\in{\mathrm{L}}^{1}([0,T];[0,+\infty))$, Equation (5.1) has a unique solution $X^{x,{\mathbf{u}}}$, given by
As in the stochastic case, $X$ describes the evolution of the production capacity of a representative agent, which depreciates at a rate $\delta$ and can be increased by choosing the investment rate $\mathbf{u}$}. Also in the deterministic case we are in an irreversible investment decisions setting, given the non-negativity of the control.
Next, we consider the discounted net profit functional
where $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. The control $\mathbf{u}$} is chosen in either of the following two classes of admissible controls: if $T<+\infty$,
@@
 $\qquad\text{ and }\int_{0}^{+\infty}{\mathrm{e}}^{-\rho s}u_{s}^{2}\,{\mathrm{d}}s<+\infty\biggr\}\,.$ (5.5)
From an economic point of view, the deterministic model captures an economy in which the production capacity of the representative firm is not affected by random shocks. Its evolution over time is, thus, uniquely determined by the investment efforts to increase it.
@@
Since the representative agent is chosen randomly by picking her/his initial production capacity level $x$ according to the initial distribution $\nu_{0}$, the assumption above ensures that, for any ${\mathbf{u}}\in{\mathcal{U}}_{T}$, $X_{s}^{x,{\mathbf{u}}}>0$, for all $s\in[0,T]$. Thus, the production capacity level of the representative agent is never negative.
@@
[…]
@@
Dimostrazione.
Fix ${\mathbf{u}}\in{\mathcal{U}}_{T}$ and $p_{0}$ as above. The fact that $p^{p_{0},{\mathbf{u}}}$ is a solution to (5.6) can be verified via standard computations (see, e.g., [7, Example 2.5.2] or [10, Lemma 4.15]). Uniqueness follows from an argument similar to the one used in the proof of [10, Lemma 4.16].
@@
[…]
@@ 5.1. The optimization problem
In the finite time horizon case, i.e. $T<+\infty$, the following result holds, which is analogous to Proposition 3.3.
Proposition 5.5.
Fix $x>0$ and $q\in{\mathcal{Q}}_{T}$, where ${\mathcal{Q}}_{T}$ is the set defined in (3.1). Then, $\widehat{\mathbf{u}}^{(T,q)}\coloneqq z^{(T,q)}\in{\mathcal{U}}_{T}$, where $z^{(T,q)}$ is the function defined in (3.5), is an optimal control for problem (5.10).
@@
Dimostrazione.
Replicating the argument of the proof of Lemma 3.1, we get that
@@
Then, the result follows using the same reasoning as in the proof of Proposition 3.3. ∎
In the infinite time horizon case, i.e. $T=+\infty$, we have the following statement, which is the deterministic counterpart of Proposition 3.6. We state it without proof.
Proposition 5.6.
Fix $x>0$ $q\in{\mathcal{Q}}_{\infty}$, where $\mathcal{Q}$$_{\infty}$ is the set defined in (3.10). Then, $\widehat{\mathbf{u}}^{(\infty,q)}\coloneqq z^{(\infty,q)}\in{\mathcal{U}}_{\infty}$, where $z^{(\infty,q)}$ is the function defined in (3.9), is an optimal control for problem (5.10). Moreover, $\widehat{\mathbf{u}}^{(\infty,q)}$ is essentially unique, i.e., if $\overline{{\mathbf{u}}}^{(\infty,q)}\in{\mathcal{U}}_{\infty}$ is an optimal control for problem (5.10) different from $\widehat{\mathbf{u}}^{(\infty,q)}$, then
@@ 5.2. Existence and uniqueness of equilibria
Also in the deterministic mean-field game previously introduced, the search for an equilibrium boils down to finding a fixed point of a suitable map. According to Definition 5.4, this map is $(q_{s})_{s\in[0,T]}\mapsto\left(\int_{0}^{+\infty}x\,m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}(s,x)\,{\mathrm{d}}x\right)_{s\in[0,T]}$, where $\widehat{\mathbf{u}}^{(T,q)}$ is the optimal control for the optimization problem (5.10) – whose expression is given either in Proposition 5.5, in the case $T<+\infty$, or in Proposition 5.6, in the case $T=+\infty$ – and $m^{m_{0},\widehat{\mathbf{u}}^{(T,q)}}$ is the unique solution to (5.6), with initial probability density $m_{0}$ and control $\widehat{\mathbf{u}}^{(T,q)}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Finally, we get the following result, which establishes existence and uniqueness of an equilibrium also for the deterministic mean-field game. We omit its proof, since it follows the same lines of the proofs of Theorems 4.9 and 4.12, except for minor adaptations.
Theorem 5.8.
[…]